# How Does Microservice Granularity Impact Energy Consumption and Performance? A Controlled Experiment

### v1

@@ VI Threats to Validity
In this section, we discuss several possible threats to validity using the dimensions from Wohlin et al. [18]. Regarding construct validity, one potential threat is our used measurement tools. We used Powerstat for energy consumption measurement, assuming its accuracy and consistency. Since it is based on RAPL, it includes both CPU and DRAM power. Preliminary tests with other tools like PowerTop showed consistently lower readings (about 800 J less) than Powerstat. Both are software-based estimation tools rather than precise measurements and cannot match the accuracy of hardware-based tools like Watts Up Pro used in similar studies [15]. While this may introduce unreliable absolute measurements, using software-based estimation tools is an accepted practice in comparative software energy research. We therefore trust that Powerstat provided consistent relative measurements across experimental treatments, allowing valid comparisons.

### v2

@@ VI Threats to Validity
In this section, we discuss several possible threats to validity using the dimensions from Wohlin2024. Regarding construct validity, one potential threat is our used measurement tools. We used Powerstat for energy consumption measurement, assuming its accuracy and consistency. Since it is based on RAPL, it includes both CPU and DRAM power. Preliminary tests with other tools like PowerTop showed consistently lower readings (about 800 J less) than Powerstat. Both are software-based estimation tools rather than precise measurements and cannot match the accuracy of hardware-based tools like Watts Up Pro used in similar studies [dinga2023empirical]. While this may introduce unreliable absolute measurements, using software-based estimation tools is an accepted practice in comparative software energy research. We therefore trust that Powerstat provided consistent relative measurements across experimental treatments, allowing valid comparisons.